The length of a rectangle is 5 ft more than the width. If the perimeter must be at least 38ft but no more than 62 ft, find the possible values for the width.
Answer by edjones(8007)

L=w+5
2L+2w=P
.
2(w+5)+2w>=38
2w+10+2w>=38
4w+10>=38
4w+10-10>=38-10
4w=28
4w/4>=28/4
w>=7
.
2(w+5)+2w<=62
2w+10+2w<=62
4w+10<=62
4w+10-10<=62-10
4w<=52
w<=13
.
7<=w<=13
.
Ed